# An Algorithmic Approach to Finding Degree-Doubling Nodes in Oriented Graphs

### v3

@@ Abstract
Seymour’s Second Neighborhood Conjecture claims that there will always exist a node whose out-degree doubles in the square of an oriented graph. This paper first presents a data structure GLOVER (Graph Level Order) which is a partitioning of the nodes into a total order of neighborhoods. This data structure allows us to construct a decreasing sequence of subsets, ensuring that no counterexample can exist. Further, this algorithmic approach allows us to know where the node is. Later findings extend the results to multiple nodes. This is true for all oriented graphs. These findings also have practical applications in network optimization and algorithm design.
@@ 4 Initial Lemmas
Case 2: Assume that $d^{+}(v_{0})=1$. Then $v_{0}$ is connected to some node $v_{1}\in G$. Since $v_{0}$ is a minimum out-degree node, we know that $d^{+}(v_{1})\geq d^{+}(v_{0})=1$. Because $G$ is an oriented graph and $(v_{0},v_{1})\in G$, it follows that $(v_{1},v_{0})\notin G$. Therefore, $v_{1}$ must be connected to another node $v_{2}\neq v_{0}$. This means that $(v_{1},v_{2})\in G$. This implies that in $G^{2}$, both $(v_{0},v_{1})$ and $(v_{0},v_{2})$ exist, giving $v_{0}$ an out-degree of at least 2, thus doubling its out-degree.
Case 3: Assume that $d^{+}(v_{0})=2$. Then $v_{0}$ is connected to two nodes, $v_{1}$ and $v_{2}\in G$. Since $v_{0}$ has the minimum out-degree, we know $d^{+}(v_{1})\geq 2$ and $d^{+}(v_{2})\geq 2$. As $G$ is an oriented graph, $(v_{1},v_{0})$ and $(v_{2},v_{0})$ are not in $G$. Consider the subgraph $G[v_{1},v_{2}]$, induced on the nodes ${v_{1},v_{2}}$. There is at most one directed edge between $v_{1}$ and $v_{2}$ in $G$, meaning one of these nodes has an out-degree of at least 2 outside $G[v_{1},v_{2}]$, connecting to other nodes. Let these nodes be $v_{3}$ and $v_{4}$. Since $G$ is oriented, $v_{0}$ cannot be one of these nodes. Hence, in $G^{2}$, $v_{0}$ will connect to $v_{1}$, $v_{2}$, $v_{3}$, and $v_{4}$, giving it an out-degree of at least 4, thus doubling its out-degree.
@@
Assume $d^{+}(v_{0})=3$, and let $v_{0}$ be connected to nodes $v_{1}$, $v_{2}$, $v_{3}\in V(G)$. Consider the subgraph $G[v_{1},v_{2},v_{3}]$, which is the subgraph induced by the set $\{v_{1},v_{2},v_{3}\}$. Now, assume that one of $v_{1},v_{2},v_{3}$, say $v_{1}$, has an out-degree of 0 within the subgraph $G[v_{1},v_{2},v_{3}]$. This implies that $v_{1}$ must connect to nodes outside of $\{v_{1},v_{2},v_{3}\}$, as $v_{1}$ has no out-neighbors within the induced subgraph.
Since $v_{0}$ is connected to $v_{1}$, and $v_{1}$ has out-neighbors outside of $\{v_{1},v_{2},v_{3}\}$, the distance-2 neighbors of $v_{0}$ in $G^{2}$ will include these additional neighbors of $v_{1}$, As a result, $v_{0}$ will have at least twice as many out-neighbors in $G^{2}$ compared to its out-degree in G.
Thus, $v_{0}$’s out-degree will at least double in $G^{2}$. ∎
@@
Assume $d^{+}(v_{0})=3,$ so $v_{0}$ is connected to nodes $v_{1},v_{2},v_{3}\in V(G).$ Since $v_{0}$ is a minimum out-degree node, we know $d^{+}(v_{i})\geq 3$ for each $v_{i}\in\{v_{1},v_{2},v_{3}\}$. Let $G[v_{1},v_{2},v_{3}]$ represent the subgraph induced by $v_{1},v_{2},v_{3}$.
We can reason that if any of $v_{1},$ $v_{2},$ or $v_{3}$ has an out-degree greater than 3, the out-degree of $v_{0}$ will at least double in $G^{2}$. Therefore, assume that all three nodes have an out-degree of exactly 3 in $G$, and by Lemma 4.2(Minimum Out-Degree 3 with Neighbor Out-Degree 0) , none of $v_{1}$,$v_{2},$ or $v_{3}$ have out-degree 0 in $G[v_{1},v_{2},v_{3}]$.
Each node $v_{1},v_{2},v_{3}$ must connect to two nodes in $G-G[v_{1},v_{2},v_{3}].$ If the set of nodes to which $v_{1},v_{2},v_{3}$ connect in $G-G[v_{1},v_{2},v_{3}]$ contains more than 2 distinct nodes, the out-degree of $v_{0}$ will at least double in $G^{2}$. Therefore, assume they connect to exactly 2 nodes outside of $G[v_{1},v_{2},v_{3}]$, say $v_{4}$ and $v_{5}$.
@@
Thus, either $v_{0}$ or one of its neighbors will have its out-degree at least double in $G^{2}$. ∎
@@ 5 Theoretical Foundation
For a node $x\in V(G)$ in an oriented graph $G$, we say that x has the decreasing sequence property if the size of its out-neighborhood is strictly greater than the size of its second out-neighborhood, i.e.,$|N^{+}(x)|>|N^{++}(x)|$,where $N^{+}(x)$ denotes the set of out-neighbors of $x$, and $N^{++}(x)$ denotes the set of out-neighbors of the nodes in $N^{+}(x)$.
@@
Let $G$ be a graph with a minimum degree node $v_{0}$, where all nodes satisfy the Decreasing Sequence Property (DSP). Construct a series of containers $C_{1},C_{2},\ldots,C_{k}$ relative to $v_{0}$, where each container $C_{i}$ consists of nodes ordered by their adjacency and connectivity properties with respect to $v_{0}$ and preceding containers. This structure imposes a total order on the nodes in $G$, where every node in $C_{i}$ is ordered before all nodes in $C_{i+1}$, establishing a ranking aligned with the DSP.
@@ 10.2 Theorem Statement
Degree Doubling by Container Density In an oriented graph $G$ with minimum degree node $v_{0}$ and containers $C_{1},C_{2},...,C_{n}$, as the containers shrink in size as their distance from the minimum degree node increases, and as the required interior degrees rise, a collision is inevitable. This collision can manifest in two forms: either the size of the container reaches a point where it can no longer support the increasing interior degrees (resulting in nodes that cannot maintain their minimum degree requirements), or smaller containers force a doubling of degrees that contradicts the capacity for interior arcs. The structure of the graph, particularly in small sizes, inherently limits these growth dynamics, leading to contradictions that must be addressed.
@@
For even values of $k$, such as $k=4$ or $k=6$, the symmetry and density required for a complete graph $K_{n}$ cannot form directly at any container level. This is because the degree-doubling effect and the odd-number dependency of $K_{2i+1}$ mean that the structural requirements for even $n$ are never satisfied in full. Instead, the structure transitions through intermediate states between odd complete graphs, and complete graphs may appear to ”exist” in a fractional or approximate form between $K_{2i-1}$ and $K_{2i+1}$.
@@
This emergence of orientations of complete graphs as containers decrease in size adds a layer of symmetry and beauty to the Seymour Second Neighborhood Conjecture. There was never a question of these arising. Some questioned multiple nodes. Some questioned zero nodes. A complete graph occurrence though is a thing of beauty and it helps give rise to another result.
@@
Identifying regularity was key to understanding why multiple nodes’ degrees can double in certain graphs. The regularity is the key factor that helps explain why multiple nodes can have their degrees double. The regularity is the central feature that ties everything together.
Degree doubling is essentially a consequence of the underlying regularity. When a graph’s structure is highly regular, particularly in terms of interior and exterior degrees, it becomes easier to predict and verify that certain structural properties (like degree doubling) will occur at specific stages.
@@ 11 Conclusion and Future Work
With a solution now in hand, its time to begin looking forward. The Seymour Second Neighborhood Conjecture can be represented as a social media problem instead of an oriented graph. Here, the nodes are people, where noone follows back. The same question is of interest, is there anyone whose degree doubles, except we know the answer is yes. We just need to implement the algorithm to find the person
We could have the graph represent epidemiology. Here the nodes be infected people, and the arcs would be one way infections. The question of interest would be can we find the person who is a super-spreader.
@@ 14 Appendix 2: Applicability of the Graph Level Order
A total order is a linear ranking. There are no containers or buckets to make extra comparisons. Everything is sorted. That may not be what we desite though. Suppose we have two metrics, like age and income. Those two metrics cannot be compared on the same scale. They’d need to be grouped by one metric and compared by the other.
The Graph Level Order (GLOVER) is able to do these multi level comparisons. GLOVER does this by the usage of containers. These containers arrange elements in sequence. Elements within each level maintain an internal order. The containers themselves are comparable across ranked layers. This structure allows for multi-level comparisons. These can happen both within a single level and between levels. This makes GLOVER a higher-level mathematical framework suitable for multi-layered, multi-dimensional data.
Just as the SSNC was proved with the GLOVER data structure, there are more proofs and algorithms that this method is capable of handling.

### v4

@@ Abstract
Seymour’s Second Neighborhood Conjecture claims that there will always exist a node whose out-degree doubles in the square of an oriented graph. This paper first presents a data structure GLOVER (Graph Level Order) which is a partitioning of the nodes into a total order of neighborhoods. This data structure allows us to introduce a total order on oriented graphs and construct a decreasing sequence of subsets. This decreasing sequence of subsets is what ensures that no counterexample can exist. Further, this algorithmic approach allows us to know where the node is. Later findings extend the results to multiple nodes. This is true for all oriented graphs. These findings also have practical applications in network optimization and algorithm design.
@@ 4 Initial Lemmas
Case 2: Assume that $d^{+}(v_{0})=1$. Then $v_{0}$ is connected to some node $v_{1}\in G$. We know that $d^{+}(v_{1})\geq d^{+}(v_{0})=1$. This node $v_{1}$ must be connected to another node $v_{2}\neq v_{0}$. This means that $(v_{1},v_{2})\in G$. This implies that in $G^{2}$, both $(v_{0},v_{1})$ and $(v_{0},v_{2})$ exist, giving $v_{0}$ an out-degree of at least 2, thus doubling its out-degree.
Case 3: Assume that $d^{+}(v_{0})=2$. Then $v_{0}$ is connected to two nodes, $v_{1}$ and $v_{2}\in G$. We know $d^{+}(v_{1})\geq 2$ and $d^{+}(v_{2})\geq 2$. Consider the induced subgraph $G[v_{1},v_{2}]$. There is at most one directed edge between $v_{1}$ and $v_{2}$ in this induced subgraph. This means that one of these nodes has an out-degree of at least 2 outside $G[v_{1},v_{2}]$, connecting to other nodes. Let these nodes be $v_{3}$ and $v_{4}$. The node $v_{0}$ cannot be one of these nodes. Hence, in $G^{2}$, $v_{0}$ will connect to $v_{1}$, $v_{2}$, $v_{3}$, and $v_{4}$, giving it an out-degree of at least 4. Thus, its out-degree will double.
@@
Assume $d^{+}(v_{0})=3$, and let $v_{0}$ be connected to nodes $v_{1}$, $v_{2}$, $v_{3}\in V(G)$. Consider the induced subgraph $G[v_{1},v_{2},v_{3}]$. Now, assume that $v_{1}$, has an out-degree of 0 within $G[v_{1},v_{2},v_{3}]$. This implies that $v_{1}$ must connect to 3 nodes outside of $\{v_{1},v_{2},v_{3}\}$.
Since $v_{0}$ is connected to $v_{1}$, and $v_{1}$ has 3 out-neighbors outside of $\{v_{1},v_{2},v_{3}\}$, the distance-2 neighbors of $v_{0}$ in $G^{2}$ will include these additional neighbors of $v_{1}$, As a result, $v_{0}$ will have at least twice as many out-neighbors in $G^{2}$ compared to its out-degree in $G$.
@@
Assume $d^{+}(v_{0})=3,$ so $v_{0}$ is connected to nodes $v_{1},v_{2},v_{3}\in V(G).$ We know that $d^{+}(v_{i})\geq 3$ for each $v_{i}\in\{v_{1},v_{2},v_{3}\}$. Let $G[v_{1},v_{2},v_{3}]$ represent the induced subgraph by $v_{1},v_{2},v_{3}$.
We can reason that if any of $v_{1},$ $v_{2},$ or $v_{3}$ has an out-degree greater than 3, the out-degree of $v_{0}$ will at least double in $G^{2}$. Therefore, assume that all three nodes have an out-degree of exactly 3 in $G$.
Each node $v_{1},v_{2},v_{3}$ must connect to two nodes in $G-G[v_{1},v_{2},v_{3}].$ If the set of nodes to which $v_{1},v_{2},v_{3}$ connect in $G-G[v_{1},v_{2},v_{3}]$ contains more than 2 distinct nodes, the out-degree of $v_{0}$ will at least double in $G^{2}$. Therefore, assume they all connect to exactly 2 nodes outside of $G[v_{1},v_{2},v_{3}]$, say $v_{4}$ and $v_{5}$.
@@ 5 Theoretical Foundation
For a node $x\in V(G)$ in an oriented graph $G$, we say that x has the decreasing sequence property if the size of its out-neighborhood is strictly greater than the size of its second out-neighborhood, i.e.,$|N^{+}(x)|>|N^{++}(x)|$.
@@
Let $G$ be a graph with a minimum degree node $v_{0}$. Further, let all nodes in $G$ satisfy the Decreasing Sequence Property (DSP). Construct a series of containers $C_{1},C_{2},\ldots,C_{k}$ ordered by distance relative to $v_{0}$. Each container $C_{i}$ will consist of nodes at that respective distance from $v_{0}$. Then this structure imposes a total order on the nodes in $G$ i.e. every node in $C_{i}$ is ordered before all nodes in $C_{i+1}$.
@@ 10.2 Theorem Statement
Degree Doubling by Container Density In an oriented graph $G$ with minimum degree node $v_{0}$ and containers $C_{1},C_{2},...,C_{n}$, as the containers shrink in size as their distance from the minimum degree node increases, and as the required interior degrees rise, a collision is inevitable.
There are two possibilities:
  1. 1.
    As the containers shrink in size and the interior degrees grow in size, we reach a point where no feasible oriented graph is possible to accommodate the required set of nodes with interior degrees.
  2. 2.
    For small minimum degree nodes $v_{0}$, The container’s shrinkage is not fast enough to cause this collision.
@@
In small graphs like 4.3 (Minimum Out-Degree 3 with Neighbors 1), we saw that this collision did not occur. To understand the reasoning behind this, understand that the container sizes are decreasing at the rate of 1, and the interior degrees are increasing at a rate of 1. The problem is that the number of arcs inside an oriented graph is growing at a rate of $O(n^{2})$. So for small graphs, we are just not going out far enough to catch up. For larger graphs though, we do start to see this.
@@
For even values of $k$, such as $k=4$ or $k=6$, the symmetry and density required for a complete graph $K_{n}$ cannot form directly at any container level. This is because the degree-doubling effect and the odd-number dependency of $K_{2i+1}$ mean that the requirements for even $n$ are never satisfied in full. Instead, it transitions through intermediate states between odd complete graphs, and complete graphs may appear to ”exist” in a fractional or approximate form between $K_{2i-1}$ and $K_{2i+1}$.
@@
This emergence of orientations of complete graphs as containers decrease in size adds a layer of beauty to the Seymour Second Neighborhood Conjecture. There was never a question of these arising. Some questioned multiple nodes. Some questioned zero nodes. A complete graph occurrence though is a thing of beauty and it helps give rise to another result.
@@ 11 Conclusion and Future Work
In this work, we explored the Seymour Second Neighborhood Conjecture in oriented graphs. We did so through the lenses of graph theory, combinatorics, algorithms and data structures. This research helped us to use many mathematical tools like both total orders, strong and weak induction, recurrence relations, and the divide and conquer paradigm.
One of the key insights of this paper was the development of an algorithm that identifies nodes whose degree must double in any oriented graph. This effectively solves the Seymour Second Neighborhood Conjecture with an answer in the affirmative. The interesting thing about this algorithm was that it did not require any assumptions about the oriented graph or where the degree doubling node was in the graph. Instead the algorithm selects a minimum degree node and is able to partition the graph into containers and form a decreasing sequence of exteriors that is eventually bound to lead us to the solution.
This approach introduced a new perspective on the Seymour Second Neighborhood Conjecture. It was no longer framed as merely a graph theory problem. By utilizing the divide and conquer paradigm and solving different parts independently, this algorithm shows that the Seymour Second Neighborhood Conjecture could effective be treated with computer science techniques. This suggests that there may be other open problems in mathematics that are not being given enough attention by computer scientists. Alternatively, it may be that there is not enough interdisciplinary communication about these problems and the potential techniques to solve them.
@@
With a solution now in hand, its time to begin looking forward. The Seymour Second Neighborhood Conjecture can be represented as a social media problem instead of an oriented graph. Here, the nodes are people, where no one follows back. Now instead of wondering if there is a person whose sphere of influence doubles, we would simply be searching for that person. We just need to implement the algorithm to find the person. We could have the graph represent epidemiology. Here the nodes be infected people, and the arcs would be one way infections. The question of interest would be can we implement the algorithm to find the person who is a super-spreader, and do so quickly? There are many more real life examples where we can apply this problem and this algorithm. There are fields like security, telecommunications, urban planning, neural networks. There are too many to name in this short paper.
@@ 14 Appendix 2: Applicability of the Graph Level Order
The deciding factor that held the algorithm together and helped solve the Seymour Second Neighborhood Conjecture was the Graph Level Order (GLOVER) data structure. This research did not begin as an investigation into data structures. Instead, a series of operations were conducted on the oriented graphs - partitioning, ordering those partitions, adding interior and exterior arcs, until what was left was an ordering of both the partition and nodes within the partition. Further, these interior and exterior arcs allow us to define relationships based on two different metrics, an outer metric and an inner metric. Unlike the Seymour Second Neighborhood Conjecture, exterior arcs do not always have to only be defined to the next container. That was a problem specific definition.
Moving forward, the GLOVER data structure should be well suited to tackle problems in graph theory and beyond. By introducing containers, another way of doing induction on graphs has been introduced. The concept of divide and conquer, which has dominated much of computer science, shows up in this paper as well. We are able to treat these containers independent of one another and solve their problems locally before bringing them back to the global problem. This shows the how versatile the GLOVER data structure is moving forward.
What makes GLOVER special are things like total order on containers. Now there is a two way ranking of nodes, as opposed to traditional one way rankings like lexicographical sorting. This agrees with many real world systems where we have dual metrics, that are often competing, like price and performance. This is hard to measure on a single scale.
Next these nodes are not just treated like clusters inside these containers, but like nodes in graph theory. The interior and exterior arcs allow for common graph theory algorithms and techniques like Prim’s algorithm, breadth first search, or centrality to be called. What is even more interesting is that because we have identified the interior and exterior arcs already, we can choose to run these algorithms only on interior arcs or only on exterior arcs.
This is still a new data structure, so we are learning a lot about it. There are many applications that have not been tried yet. However, since it brings the strengths of a total order and a graph into one unified structure, it should be able to model complex relationships. This should include domains of machine learning like clustering, dynamic scheduling, resource management, project management, and natural language processing.